OpenStudy (anonymous):

y=(x/k)+k where k is constant tangent 4y=x^2, i) find k, ii) find P

OpenStudy (freckles):

Are you wanting to find the tangent line to 4y=x^2?

OpenStudy (freckles):

You will need to first differentiate 4y=x^2?

OpenStudy (anonymous):

I need to find the value of constant K using the information above, i meant 4y=x^2 is the tangent to the curve y=(x/k)+k and joins the curve at point P, finding point P is the second part of the question where i just have to do simultaneous equations im pretty sure

OpenStudy (freckles):

Well I still think we need to differentiate 4y=x^2 to find the general slope formula and compare it to the slope of the line which is 1/k

OpenStudy (anonymous):

okay thanks

OpenStudy (freckles):

Well have you done that yet?

OpenStudy (anonymous):

yup

OpenStudy (freckles):

which is?

OpenStudy (anonymous):

x/2

OpenStudy (freckles):

Ok so say we call the point that we want to find is (a,b)

OpenStudy (freckles):

so the derivative evaluated at x=a is a/2

OpenStudy (freckles):

So we have a/2=1/k

OpenStudy (freckles):

So now solve that for k

OpenStudy (freckles):

just a picture if you think it would be helpful